OpenStudy (swissgirl):

Use the one to one correspondences ln: (0, infinity) --> R f: ( 2, infinity) --> (0, infinity) Where f(x)=x-2 to describe a one to one correspondence from (0,infinity) onto (2, infinity)

OpenStudy (swissgirl):

Umm I so dont follow what to do here

OpenStudy (anonymous):

on first glace it looks like "first take the log, then subtract 2"

OpenStudy (anonymous):

hmm maybe "add 2"?

OpenStudy (swissgirl):

idk

OpenStudy (swissgirl):

like i am a lil stimped

OpenStudy (swissgirl):

stumped*

OpenStudy (anonymous):

yeah this is weird if you want a one to one corresponedence between \((2,\infty)\) and \((0,\infty)\) what is wrong with \(x:\to x-2\)

OpenStudy (swissgirl):

Nothing -_-

OpenStudy (swissgirl):

I mean is that all that needs to be done? This is nerdy

OpenStudy (swissgirl):

Umm how about from (2, infinity) onto R x: ln(x-2)?

OpenStudy (anonymous):

that would certainly do it

OpenStudy (swissgirl):

yayyyy, Its weird that it is soo simple but i guess i wont complain

OpenStudy (anonymous):

sometimes it is so simple it seems hard

OpenStudy (swissgirl):

Now from R onto (2, infinity) wld it be e^(x-2)?

OpenStudy (swissgirl):

@nbouscal is this correct?

OpenStudy (swissgirl):

@mukushla

OpenStudy (anonymous):

lol....idk...

OpenStudy (swissgirl):

thanks :)

OpenStudy (swissgirl):

@satellite73 can u check this?

OpenStudy (kinggeorge):

Here's what I would do. Find a 1-1 correspondence (a bijection) from \((2,\infty)\) to \((0,\infty)\). Then take the inverse.

OpenStudy (swissgirl):

umm so i did the other examples incorrect too?

OpenStudy (kinggeorge):

Let me read over this more carefully.

OpenStudy (kinggeorge):

So satellite went from (2, infinity) to (0, infinity) at first. We need to go the other way. It's not much harder. You just need to use \(f^{-1}=x+2\).

OpenStudy (kinggeorge):

From (2, infinity) to R, you nailed it. It's just \(\ln(f(x))=\ln(x-2)\).

OpenStudy (kinggeorge):

From R to (2, infinity), it's the inverse of what I just wrote above. \[f^{-1}(\ln^{-1}(x))=f^{-1}(e^x)=e^x+2\]

OpenStudy (swissgirl):

Alrrighhtttttt. YAYYYYYYY

OpenStudy (swissgirl):

At

OpenStudy (swissgirl):

least i accomplished smth this evening

OpenStudy (swissgirl):

Thanks :)

OpenStudy (kinggeorge):

You're welcome.